A stone ot mass 2.0 kg to pulled a long a horizontal path by a rope which makes an angel 60° with the horizontal neglecting frictions. Calculate the tension in the rope of the body accelerate uniformly at 5.0 m5-2


1
Expert's answer
2021-06-20T18:04:06-0400

Gives

Mass(m)=2.0kg

a=5m/sec2

"\theta=60°"

"F=ma"

"F=2\times5=10N"

Tansion horizontal components

F=Tcos"\theta"

"T=\frac{F}{cos\theta}"

"T=\frac{10}{cos60°}=20N"